OpenStudy (anonymous):

An ice skater is traveling in a straight, horizontal line on the ice with a velocity of v = 8.5 m/s in the positive x-direction. The coefficient of kinetic friction between the skates and the ice is μk = 0.19. How long, th in seconds, will it take for the skater's velocity to drop by half?

OpenStudy (farcher):

Acceleration a is - 0.19 g m/s^2 and use v = u + at. u = 8.5 m/s and v = 8.5/2 m/s.

OpenStudy (anonymous):

thank you so much